OpenStudy (anonymous):

use the associative property of multiplication to write an equivalent expression. 8 (np)

OpenStudy (kropot72):

\[(\sqrt{8}n)\times(\sqrt{8}p)\]

OpenStudy (kropot72):

Do you understand my result?